OpenStudy (anonymous):

At 1 p.m., Lin records an outside temperature of –14.8°F. At 3 p.m., she records the outside temperature and notices the temperature had increased by 22.6 degrees. What was the temperature that she recorded at 3 p.m.?

OpenStudy (anonymous):

I think its 7.8

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

is it 37.4

OpenStudy (agl202):

You will want to add -14.8 and 22.6 and you will get 7.8 so that will be your answer. 7.8 degrees.

OpenStudy (anonymous):

ok just trying to make sure thx

OpenStudy (anonymous):

The variable t represents the temperature at 3 p.m. Which equation would best model this word problem? The temperature at the summit of Mt. Mitchell dropped 23°F between 3 p.m. and 9 p.m. The temperature at 9 p.m. was –10°F. What was the temperature at 3 p.m.? A. t – 23 = –10 B. 10 – 23 = t C. 10 + 23 = t D. 23 – t = –10

OpenStudy (anonymous):

Im thinking D but ins

OpenStudy (anonymous):

or maybe B.

OpenStudy (agl202):

No, I think ur right with D. One question per post plz.

OpenStudy (anonymous):

ok thx